Given the Total cost Function, TC = 5,600 + 27Q - 6Q² - .5Q³, solve for AC, MC, AVC and AFC





1
Expert's answer
2021-11-03T10:40:58-0400

"Given;\\

TC=5600+27Q−6Q^2−0.5Q^3\\

AC=\frac{TC}{Q}\\ =\frac{5600}{Q}+27−6Q−0.5Q^2\\

MC=\frac{dTC}{dQ}\\ =27−12Q−1.5Q^2\\

AVC=\frac{VC}{Q}\\ =\frac{27Q−6Q^2−0.5Q^3}{Q}\\ =27−6Q−0.5Q^2\\

AFC=\frac{FC}{Q}\\ =\frac{5600}{Q}"